# Do Robot Snakes Dream like Electric Sheep? Investigating the Effects of Architectural Inductive Biases on Hallucination

### v1

@@ Abstract
The growth in prominence of large language models (LLMs) in everyday life can be largely attributed to their generative abilities, yet some of this is also owed to the risks and costs associated with their use. On one front is their tendency to hallucinate false or misleading information, limiting their reliability. On another is the increasing focus on the computational limitations associated with traditional self-attention based LLMs, which has brought about new alternatives, in particular recurrent models, meant to overcome them. Yet it remains uncommon to consider these two concerns simultaneously. Do changes in architecture exacerbate/alleviate existing concerns about hallucinations? Do they affect how and where they occur? Through an extensive evaluation, we study how these architecture-based inductive biases affect the propensity to hallucinate. While hallucination remains a general phenomenon not limited to specific architectures, the situations in which they occur and the ease with which specific types of hallucinations can be induced can significantly differ based on the model architecture. These findings highlight the need for better understanding both these problems in conjunction with each other, as well as consider how to design more universal techniques for handling hallucinations.
@@ 9 Acknowledgements
Jerry Huang received partial support under the forms of a National Science and Engineering Research Council (NSERC) Canada Graduate Scholarship, a Fonds de Recherche du Québec Nature et Technologies (FRQNT) Training Scholarship, and a Hydro-Québec Excellence Scholarship over the course of this work. Sarath Chandar is supported by a Canada CIFAR AI Chair, the Canada Research Chair in Lifelong Machine Learning, and a NSERC Discovery Grant.
@@ B.1 Numerical Results
Numerical results are provided in the following tables
  • Figure 4 and Table 3 contains results for all individual tasks.
@@
Hallu. Detection Instr. Following QA Reading Comp. Sum. Fact-Checking

### v2

@@ Abstract
The growth in prominence of large language models (LLMs) in everyday life can be largely attributed to their generative abilities, but some of this is also due to the risks and costs associated with their use. On one front is their tendency to hallucinate false or misleading information, limiting their reliability. On another is the increasing focus on the computational limitations associated with traditional self-attention based LLMs, which has brought about new alternatives, in particular recurrent models, meant to overcome them. However, it remains rare to consider these two concerns simultaneously. Do changes in architecture exacerbate/alleviate existing concerns about hallucinations? Do they affect how and where they occur? Through an extensive evaluation, we study how these architecture-based inductive biases affect the propensity to hallucinate. Although hallucination remains a general phenomenon not limited to specific architectures, the situations in which they occur and the ease with which specific types of hallucinations can be induced can differ significantly depending on the model architecture. These findings highlight the need for better understanding both these problems in conjunction with each other, and consider how to design more universal techniques for handling hallucinations.
@@ 9 Acknowledgements
Jerry Huang received partial support in the form of a National Science and Engineering Research Council (NSERC) Canada Graduate Scholarship, a Fonds de Recherche du Québec Nature et Technologies (FRQNT) Training Scholarship, and a Hydro-Québec Excellence Scholarship during the course of this work. Sarath Chandar is supported by a Canada CIFAR AI Chair, the Canada Research Chair in Lifelong Machine Learning, and a NSERC Discovery Grant.
@@ B.1 Numerical Results
Numerical results are provided in the following tables:
  • Figure 4 and Table 3 contain results for all individual tasks.
@@
Model Hallucination Instruction Closed-Book Reading Summarization Fact-Checking
Detection Following Open-Domain QA Comprehension